Using zero coupon bonds You know the information on three bonds that make annual coupon payments, if any. Bond Maturity 1) Whplz show steps
Answer #1

Price of ZCB= Amount/ (1+r)n where r is Yield to maturity rate , n is periods, amount = face value

1) Hence price of A = 1000/(1.05)1= $952.38

Price of bond B = 1000/(1.0585)2=$944.73

2)The zero coupon bond yield of 1 year bond is 5% and zero coupon yield of 2 year bond is 5.85%

3) and 4) cant be answered as neither the Yield to maturity is given nor the price is given of Bond C